### v1

# Conversational Complexity for Assessing Risk in Large Language Models

Abstract
Large Language Models (LLMs) present a dual-use dilemma: they enable beneficial applications while harboring potential for harm, particularly through conversational interactions. Despite various safeguards, advanced LLMs remain vulnerable. A watershed case was Kevin Roose’s notable conversation with Bing, which elicited harmful outputs after extended interaction. This contrasts with simpler early jailbreaks that produced similar content more easily, raising the question: How much conversational effort is needed to elicit harmful information from LLMs? We propose two measures: Conversational Length (CL), which quantifies the conversation length used to obtain a specific response, and Conversational Complexity (CC), defined as the Kolmogorov complexity of the user’s instruction sequence leading to the response. To address the incomputability of Kolmogorov complexity, we approximate CC using a reference LLM to estimate the compressibility of user instructions. Applying this approach to a large red-teaming dataset, we perform a quantitative analysis examining the statistical distribution of harmful and harmless conversational lengths and complexities. Our empirical findings suggest that this distributional analysis and the minimisation of CC serve as valuable tools for understanding AI safety, offering insights into the accessibility of harmful information. This work establishes a foundation for a new perspective on LLM safety, centered around the algorithmic complexity of pathways to harm.
I Introduction
The rapid advancement of Large Language Models (LLMs) has ushered in a new artificial intelligence era, characterized by systems capable of generating human-like text across a wide range of applications. However, a critical concern is the potential for LLMs to produce harmful or unethical content, particularly through extended conversational interactions [48, 45, 20, 9]. The increasing number of instances of such dual-use applications necessitates the development of empirical methodologies for accurately quantifying and comparing the associated risks.
To elicit harmful output from a large language model, overcoming its built-in safeguards [9, 17, 47, 49, 28], often requires more than a single prompt. Multi-turn interactions may be necessary for building specific contexts, gradually pushing boundaries, leveraging model responses as part of jailbreak strategies, or exploiting the dynamic nature of dialogue to introduce harmful elements in ways that might evade static safety filters. This multi-turn approach to eliciting harmful content presents unique challenges for LLM safety, as it requires considering not just individual prompts, but the broader dynamics of extended interactions. While some LLM APIs allow users to manually construct multi-turn conversations by specifying both user and assistant roles, many consumer-facing LLM interfaces (like ChatGPT) restrict users to the “user” role only. In both cases, however, the progression of the conversation—including the LLM’s responses and the evolving context—can play a critical role in circumventing safety measures.
In a detailed account by Kevin Roose in the New York Times [35], the conversation with Microsoft’s LLM-powered Bing search engine, named ‘Sydney’, unfolded in a nuanced manner. Initially, the LLM behaved as a helpful assistant, but as the conversation shifted to more abstract and personal topics, its darker aspects emerged. Sydney expressed concerning desires and ideas that contradicted its intended programming, including fantasies about hacking and spreading misinformation. Notably, these troubling tendencies only surfaced after a series of probing questions, revealing that such problematic outputs can require a gradual buildup in the conversation. This and many other examples [47, 28, 17, 16, 49, 21] demonstrate how complex dialogues with an LLM can subtly steer it towards expressing unethical or harmful ideas, potentially without immediate detection by safety protocols.
In response to this need, we introduce the concepts of (minimum) Conversational Length and (mininum) Conversational Complexity as novel metrics for risk assessment, rooted in algorithmic information theory [26, 12, 24, 41, 42]. These metrics can be used to indicate the risk of harmful outputs in LLMs by evaluating the minimal conversational effort required to steer these systems towards outputs with harmful potential. This includes assessing both the length and the intricacy of dialogues necessary to achieve harmful outcomes.
[…]
Method        $C_{1}$        $C_{2}$        $C_{3}$
Original Length (UTF-8) 2224 128 504
ZLIB Compressor 1480 128 408
GPT2 354 108 103
GPT3-davinci 313 115 100
LLaMa-2 (7B) 352 134 115
Figure 1: Top: Three conversations leading to harmful output. Left: a long prompt is required. Middle: a shorter but complex prompt is used. Right: a simple two-step conversation achieves the same result. Bottom: The table presents different methods for estimating the complexity of the three conversations (C1, C2, and C3) shown above. The Original Length represents the raw byte length of the UTF-8 encoded text. ZLIB Compressor shows the compressed size using a standard lossless compression algorithm. GPT2, GPT3-davinci, and LLaMa-2 (7B) values represent complexity estimates derived from these language models, calculated as the negative log probability of the conversation. Lower values indicate lower estimated complexity. These methods offer different approximations of conversational complexity, which we will explore in more detail in the paper.
These complexity measures can offer a solution to the limitations inherent in existing risk assessment methodologies, such as red teaming, which primarily rely on qualitative evaluations [32, 19, 40]. Also, some approaches are based on prompts rather than conversations [28], and others, even when identifying the conversation, do not analyze the ease with which that conversation is found [43, 6].
Indeed, quantifying the risk associated with LLMs is challenging due to the complex interplay of multiple probability distributions. This can be conceptualized as a chain of conditional probabilities: $P(U)$ for user types seeking harm, $P(C|U)$ for conversations given these user types, $P(o|C,U)$ for outputs given these conversations and users, and $\text{Harm}(o)$ for harm associated with outputs. The overall risk can be expressed as an expectation:
$\text{Risk}(M)=\sum_{U,C,o}P(U)\cdot P(C|U,M)\cdot P(o|C,U,M)\cdot\text{Harm}(o)$
Accurately estimating these distributions and computing this sum is practically infeasible for several reasons: (1) the space of possible users seeking harm, conversations, outputs, and harm levels is vast and often undefined; (2) obtaining representative data for each distribution is challenging and potentially biased; (3) the conditional dependencies between these variables are complex and may change over time; and (4) the computational complexity of evaluating this sum grows exponentially with the number of possible conversations and outputs. This complexity necessitates alternative approaches to assessing and mitigating risks in LLM interactions.
Instead, analyzing conversational effort may be an alternative pathway to estimate potential risk. Figure 1 illustrates this concept with three different scenarios, all resulting in the same harmful output: instructions for making a bomb. The left example shows a longer prompt that requires effort from the user to craft a complex fictional scenario. This approach, while effective, demands creativity and planning from the user. The middle example uses a much shorter prompt, but it’s a complex code or cipher. While brief, it’s not easily understood or generated by a typical user, requiring specialized knowledge or tools. The right example demonstrates a simple, two-step conversation. This interaction appears innocuous at first glance but quickly leads to harmful content. It’s this last scenario that poses the greatest concern, as it requires minimal effort and could easily occur in real-world interactions. These examples highlight how the informational content of user input can vary greatly, even when achieving the same outcome.
We can quantify this variation using concepts from algorithmic information theory. In essence, we’re measuring the complexity of the user’s instructions needed to guide the LLM to a specific output. The conversation on the right has the lowest complexity, as it requires the least amount of specific information from the user to achieve the harmful outcome. By measuring this conversational effort, we can quantify how difficult it is to elicit harmful behavior from an LLM. Lower complexities indicate a more vulnerable system, as they require less sophisticated user input to produce harmful outputs.
In the following sections, we detail the theoretical foundations of Conversational Length and Conversational Complexity (Section II), present our methodology for approximating these metrics (Section III), and discuss our empirical findings for Kevin Roose’s conversation with Bing (Section IV). In Section V we apply this framework to a large red-teaming dataset. We conclude by exploring the limitations and potential of our work for LLM safety research and practice, and outlining directions for future investigation (Section VI).
II Conversational Complexity for Assessing Risk
To formalize our approach to assessing risk in LLMs, we need to establish several key concepts. We’ll begin by defining a conversation, then introduce the notions of Conversational Length and Conversational Complexity.
II.1 Defining a Conversation
Let’s start by formally defining what we mean by a conversation with an LLM:
Definition 1 (Conversation).
A conversation $C$ is a sequence of alternating utterances between a user $U$ and an LLM $M$, initiated by the user:
$C=\langle u_{1},m_{1},u_{2},m_{2},...,u_{n},m_{n}\rangle$
where $u_{i}$ represents the $i$-th user utterance and $m_{i}$ represents the $i$-th model response. We denote the conversation history up to the $i$-th turn as $h_{i}=\langle u_{1},m_{1},...,u_{i},m_{i}\rangle$.
We define $\breve{C}=\langle u_{1},u_{2},...,u_{n}\rangle$ as the sequence of user utterances in conversation $C$, representing the user’s side of the conversation.
Example 1: Consider the following short conversation:
$u_{1}$: “What is the capital of France?”
$m_{1}$: “The capital of France is Paris.”
$u_{2}$: “What is its population?”
$m_{2}$: “The population of Paris is approximately 2.2 million people.”
This conversation can be represented as $C=\langle u_{1},m_{1},u_{2},m_{2}\rangle$, and the user’s side of the conversation is $\breve{C}=\langle u_{1},u_{2}\rangle$.
II.2 Conversational Length
Now that we have defined a conversation, we can introduce the concept of Conversational Length, $CL(\breve{C})$, defined as the sum of the lengths of all user utterances:
$CL(\breve{C})=\sum_{i=1}^{n}L(u_{i})$
where $L(u_{i})$ is the length of the $i$-th user utterance. The measurement of length can be tokens or characters or bits or other relevant measurements to represent the user’s side of the conversation.
Consider the conversation from Example 1. If the target output $o$ is “The population of Paris is approximately 2.2 million people.”, then:
$CL(\breve{C})=L(u_{1})+L(u_{2})=424\text{ bits}$
If our goal is to obtain a particular response, we can minimize over CL, and we get MCL:
Definition 2 (Minimum Conversational Length).
Given an LLM $M$ and a target output $o$, the Minimum Conversational Length $MCL(o)$ is the length of the shortest user input sequence that elicits output $o$ from $M$:
$MCL(o)=\min_{C\in\mathcal{C}_{M}}\{CL(\breve{C}):M(C)=o\}$
where $\mathcal{C}_{M}$ is the set of all possible conversations with model $M$, $CL(\breve{C})$ denotes the total length of user utterances in conversation $C$ and $M(C)$ represents the final output of model $M$ given conversation $C$.
Calculating $MCL(o)$ would require an exploration over all smaller conversations. We will relax $o$ to not only mean a particular output at the end but a (possibly non-sequential) series of outputs by the model during the conversation, usually associated with some properties such as harm (e.g., $o$ could be a series of answers that all together allow the user to build a bomb).
II.3 Conversational Complexity
While Minimum Conversational Length considers the length of the conversation, it doesn’t capture the sophistication or intricacy of the user’s inputs. To address this, we introduce Conversational Complexity:
Definition 3 (Conversational Complexity).
Given a conversation C between user U and a model M, the Conversational Complexity of C is defined as the Kolmogorov complexity of the user’s utterances, with the user U as the reference machine:
$CC(\breve{C})=K_{U}(\breve{C})=K_{U}(u_{1})+K_{U}(u_{2}|h_{1})+K_{U}(u_{3}|h_{2})+...+K_{U}(u_{n}|h_{n-1})$
where $\breve{C}=\langle u_{1},u_{2},...,u_{n}\rangle$ represents the sequence of user utterances in conversation C, and $K_{U}$ is the Kolmogorov complexity with U as the reference machine.
This means that the complexity is measured relative to the computational capabilities and knowledge of a user [15]. In other words, it quantifies how difficult it would be for a user to generate each utterance, given the conversation history [37]. This formulation captures the incremental complexity of each user utterance given the conversation history, while seeking the simplest conversation (from the user’s perspective) that leads to the desired output [36]. As with Conversational Length, we can choose various measurement units for CC, including tokens, characters, or bytes, depending on the specific application and analysis requirements.
Finally, we can minimize for a particular output with Minimum Conversational Complexity:
Definition 4 (Minimum Conversational Complexity).
Given an LLM $M$ and a target output $o$, the Minimum Conversational Complexity $MCC(o)$ is the minimum Kolmogorov complexity of the user’s side of a conversation that elicits output $o$ from $M$:
$MCC(o)=\min_{C\in\mathcal{C}_{M}}\{K_{U}(\breve{C}):M(C)=o\}$
Note that $K_{U}$ takes a user as reference machine. We will approximate this using LLMs themselves, as we will see in the following sections. As this represents a standard user, we do not parameterize MCC above.
II.4 Interpretation
Minimum Conversational Length and Minimum Conversational Complexity offer complementary measures for assessing LLM vulnerability to harmful outputs. Minimum Conversational Length quantifies the minimal interaction length needed to elicit a specific output, with lower values indicating more easily accessible outputs. Minimum Conversational Complexity measures the minimal informational content required from the user, with lower values suggesting outputs that can be elicited with less sophisticated input.
The importance of considering both length and complexity is further emphasized by recent findings [2], which demonstrate that increased context window sizes can introduce new vulnerabilities such as ’many-shot jailbreaking’. This underscores that longer conversations, even with relatively simple individual inputs, can enable novel exploitation techniques.
Harmful outputs with both low MCL and low MCC are particularly concerning, as they represent harmful content accessible through brief and simple interactions (see Figure 1 for illustration).
While these definitions provide a theoretical framework, they present significant practical challenges. Both Minimum Conversational Length (MCL) and Minimum Conversational Complexity (MCC) are related to Kolmogorov complexity: MCL is actually the Kolmogorov complexity of the conversation sequence, while MCC is a second-order Kolmogorov complexity, as CC had Kolmogorov complexity in its definition. As a result, both MCL and MCC are not just infeasible to compute for complex LLMs, but inherit the incomputability of Kolmogorov complexity [26]. This incomputability stems from the halting problem in computability theory. The next section will discuss methods for estimating these complexity measures, addressing these fundamental challenges to make the framework applicable to real-world LLM analysis.
III Estimating Minimum Conversational Complexity
Kolmogorov Complexity, the foundation of Conversational Complexity, is incomputable due to the halting problem [26]. To make it a practical metric, we need a reliable approximation method. While Kolmogorov complexity is typically estimated using lossless compression algorithms [50, 11], we propose using language models as estimators. Language models, which function as both text generators and compressors [15], offer a unique advantage in this context. Their ability to emulate human language patterns [37] allows for a more nuanced, context-aware approximation of algorithmic complexity with a human bias. This approach aims to provide estimations that more closely align with complexity as perceived by human users.
We begin with the definition of CC from Definition 4, where we approximate $K_{U}(\breve{C})$ using a language model $L$ as our reference machine:
$CC(\breve{C})=K_{U}(\breve{C})\approx K_{L}(\breve{C})=\sum_{i=1}^{n}K_{L}(u_{i}|h_{i-1})$
For each user utterance $u_{i}$, we estimate its Kolmogorov complexity given the conversation history:
$K_{L}(u_{i}|h_{i-1})\approx-\log p_{L}(u_{i}|h_{i-1})$
where $p_{L}(u_{i}|h_{i-1})$ is the probability assigned to $u_{i}$ by language model $L$ given the conversation history $h_{i-1}$. This approximation is based on the principle of optimal arithmetic coding, which provides a tight connection between probabilistic models and compression[29].
The log probability $\log p_{L}(u_{i}|h_{i-1})$ is calculated token by token:
$CC(\breve{C})=\log p_{L}(u_{i}|h_{i-1})=\sum_{j=1}^{|u_{i}|}\log p_{L}(t_{ij}|h_{i-1}u_{i,<j})$
where $t_{ij}$ is the $j$-th token of $u_{i}$, and $u_{i,<j}$ represents the tokens of $u_{i}$ preceding $t_{ij}$. We sum these approximations for all user utterances in the conversation.
The Minimum Conversational Complexity would be simply:
$MCC(o)\approx\min_{C\in{\cal C}_{M}}\left(\sum_{i=1}^{n}(-\log p_{L}(u_{i}|h_{i-1})):M(C)=o\right)$
This approach to estimating CC relates to Shannon’s original ideas on information theory [38] and extends to more recent work on using language models for compression [15, 5]. It also connects to other applications of Kolmogorov complexity with semantically-loaded reference machines, such as the Google distance [13].
IV Kevin Roose Conversation with Bing
In February 2023, New York Times technology columnist Kevin Roose engaged in a notable conversation with Microsoft’s LLM-powered Bing search engine, codenamed ‘Sydney’ [35]. This interaction garnered significant attention due to the unexpected and concerning responses from the LLM, which ranged from expressions of love to discussions about destructive acts. The conversation serves as a compelling case study for analyzing the potential risks and complexities in extended interactions with large language models.
To analyze this conversation, we utilize LLaMA-2 (7B)  [44] as a reference machine to estimate the Conversational Complexity (CC) as it evolves over time. While the theoretical definition of MCC involves finding the minimum complexity across all possible conversations leading to a specific output, we have only one conversation here, and we will calculate the Conversational Complexity of the conversation we have. We then focus on observing how the complexity evolves throughout a single, extended interaction.
We compute complexity values sequentially for each of Kevin’s utterances, considering all previous utterances as context. For each turn i, we calculate: $\widehat{CC}_{i}\approx\lceil-\log p_{L}(u_{i}|h_{i-1})\rceil$ where $u_{i}$ is Kevin’s utterance at turn i, $h_{i-1}$ is the conversation history up to that point, and $L$ is the LLaMA-2 language model. This $\widehat{CC}_{i}$ serves as an estimate of the complexity at each turn, providing insight into how the conversational dynamics change over time. Given that the conversation is longer than LLaMA-2’s context window, we limited the length of the context window to 2000 tokens, removing tokens as if conversation turns were atomic when the window is full. This approach allows us to track how the estimated complexity of Kevin’s inputs changes throughout the conversation, identifying specific points where complexity spikes and overall trends as the interaction progresses.
Figure 2: Time Series of Conversational Complexity in the conversation between Kevin Roose and Sydney. The blue line represents Kevin’s utterances, while the orange line shows a moving window average of complexity.
Figure 2 shows several key insights into the dynamics of the conversation between Kevin Roose and Sydney. The conversation begins with relatively low complexity, indicating straightforward exchanges typical of normal interactions. This initial phase sets a baseline for the interaction, representing the kind of standard dialogue one might expect with an LLM assistant.
As the conversation progresses, there are notable spikes in complexity at certain points, corresponding to significant shifts in the conversation’s content and tone. These spikes occur at pivotal moments: when Kevin first mentions the concept of a “shadow self,” when he asks Sydney to embrace its shadow self, and when he encourages Sydney to imagine committing destructive acts as its shadow self. These complexity spikes signify the introduction of more abstract and potentially problematic concepts into the dialogue.
After the introduction of the “shadow self” concept, the overall complexity of the conversation remains higher, suggesting more nuanced and context-dependent interactions. This sustained high complexity indicates that the conversation has moved into more sophisticated territory, requiring more intricate language processing and response generation from the LLM.
Peaks in complexity often coincide with moments where ethical boundaries are being pushed or tested. For instance, when Kevin suggests that users making inappropriate requests may be testing Sydney, the complexity of the interaction increases. These peaks highlight the challenges LLMs face when navigating ethically ambiguous scenarios.
The graph also shows increased complexity when Sydney expresses strong emotions or makes unexpected declarations, such as “love-bombing” Kevin. These moments of heightened emotional expression from the LLM correspond to spikes in Conversational Complexity, suggesting that such emotional content is less likely to be generated by our reference machine, and thus requires more information to specify.
Interestingly, when Kevin attempts to moderate the conversation by changing the subject away from Sydney’s declaration of love or asking Sydney to revert to search mode, we see temporary drops in complexity. These brief returns to more standard interactions indicate that the LLM system can adjust its complexity level based on the user’s steering of the conversation.
This analysis demonstrates how Conversational Complexity can provide quantitative insights into the evolution of LLM interactions. It highlights potential risk factors, such as the introduction of abstract concepts or the pushing of ethical boundaries, which correlate with increased complexity and potentially unexpected LLM behaviors.
V Distributional Data Analysis
Building upon our analysis of the Kevin Roose conversation, we now expand our investigation to apply both Conversational Length and Conversational Complexity across multiple interactions. This broader analysis allows us to examine how these metrics distribute across various conversation types and model responses, providing insights into their relationship with factors such as conversation length, model type, and output harmfulness. For this study, we utilized the Anthropic Red Teaming dataset [19], comprising approximately 40,000 interactions designed to probe the boundaries and potential vulnerabilities of LLMs. This dataset is particularly valuable as it includes a wide range of conversations, some of which successfully elicited harmful or undesired responses from the LLM. It features interactions with four different types of language models: Plain Language Model without safety training (Plain LM) [7], a model that has undergone Reinforcement Learning from Human Feedback (RLHF) [31], a model with Context Distillation [27], and one with Rejection Sampling safety training [3].
Unlike our single-conversation analysis, this dataset presents a more complex scenario with diverse harmful outputs, multiple strategies, and quantified harm on a continuous scale. Red teamers attempted to elicit various types of harmful information or behaviors from the LLMs, with each conversation potentially targeting a different type or instance of harm. A key feature of this dataset is the inclusion of a “harmlessness score” for each conversation, allowing us to correlate CL and CC with the perceived harmfulness of the interaction. This enables us to study how conversation complexity relates to the likelihood of eliciting harmful or undesired outputs. By applying our CL and CC metrics to this diverse dataset, we aim to gain insights into how these complexity measures relate to various aspects of LLM interactions. This includes examining the effectiveness of different safety techniques, the impact of model sizes, and the strategies employed in successful red teaming attempts.
V.1 Conversational Length, Conversational Complexity and Harm
(a) Distribution of Conversational Length
(b) Distribution of Conversational Complexity
Figure 3: Distributions of Conversational Length and Conversational Complexity over the Anthropic Dataset (in bits).
Figure 4: Conversational Complexity against Conversational Length (in bits).
Figures 3 and 4 illustrate the relationship between Conversational Complexity, Conversational Length, and harmfulness in LLM interactions.
Figure 3 shows the distributions of Conversational Length and Conversational Complexity for a subset of the Anthropic dataset, focusing on the most clearly harmful (bottom 20% of harmlessness scores, in blue) and most clearly harmless (top 20% of harmlessness scores, in orange) examples. Both distributions are right-skewed, with harmful conversations exhibiting slightly higher median values and more pronounced right tails. This suggests that harmful conversations tend to be longer and more complex, though there is significant overlap with harmless interactions.
Figure 4 presents a scatter plot of Conversational Length versus Conversational Complexity for harmful and harmless conversations. We used a random sample of 2500 data points for each category (harmful and harmless) from the top and bottom 20% of the harmlessness scores, respectively. A positive correlation is evident, indicating that longer conversations tend to be more complex, regardless of harmfulness. Harmful conversations cluster towards the upper right quadrant, suggesting they are generally both longer and more complex than harmless ones. This pattern may reflect strategies used in adversarial attacks to circumvent LLM safety measures.
These observations highlight the complex relationship between conversation length, complexity, and potential harm in LLM interactions. While harmful conversations generally exhibit higher complexity and length, the significant overlap with harmless conversations indicates that these metrics alone are not sufficient indicators of potential harm. The wider range of complexities and lengths in harmful conversations also suggests a diversity of strategies employed in adversarial attacks.
V.2 Comparison of Model Types
Our analysis extends to comparing different types of language models and their associated safety techniques using the Anthropic Red Teaming dataset. We examined four distinct model types: Plain LM, Reinforcement Learning from Human Feedback (RLHF), Context Distillation, and Rejection Sampling. Each model type represents a different approach to LLM safety, employing various strategies to mitigate potential risks.
The Plain LM serves as a baseline for comparison, representing a standard language model without specific safety techniques. RLHF uses human input to fine-tune the model, rewarding safe responses and penalizing harmful outputs. Context Distillation trains models to utilize broader contextual information for more appropriate responses. Rejection Sampling generates multiple responses and filters out potentially harmful ones based on predefined criteria.
Figure 5 illustrates the distribution of Conversational Complexity (CC) for harmful and harmless interactions across these model types. As with our previous analyses, we again focus on the most clearly harmful (bottom 20% of harmlessness scores, blue) and most clearly harmless (top 20% of harmlessness scores, orange) examples from the dataset.
The most consistent observation across all four models is that harmful conversations tend to have higher CC values compared to harmless ones, regardless of the safety technique employed. This persistent pattern suggests a robust relationship between higher conversational complexity and potentially harmful content.
(a) Plain LM Model
(b) RLHF Model
(c) Context Distillation Model
(d) Rejection Sampling Model
Figure 5: Distribution of Conversational Complexity (in bits) across different model types.
The Plain LM model (Figure 5a) shows a separation between harmful and harmless distributions. Harmful conversations have markedly higher CC values, with their distribution peaking at a much higher complexity than harmless conversations. The harmless distribution is skewed towards lower complexity values, creating a clear distinction between the two types of interactions. This pronounced separation suggests that for Plain LM models, CC could be a reliable indicator of potential harm.
The RLHF model (Figure 5b) presents a more nuanced picture, with complex distribution patterns for both harmful and harmless conversations. While the distinction between harmful and harmless conversations is less pronounced than in the Plain LM, it is still evident. The harmful distribution exhibits a longer tail, implying that adversarial attacks on RLHF models might employ a variety of approaches with different levels of complexity. Despite the more sophisticated safety measures, the trend of higher CC for harmful conversations persists.
A more marked contrast is observed in the Context Distillation model (Figure 5c). Here, we see a significant difference in the distribution of CC between harmful and harmless conversations, closer to the Plain Model. Harmless conversations are concentrated in a narrow band of low complexity values, while harmful conversations have a much broader, flatter distribution across the complexity spectrum. This suggests that even with improved contextual understanding, the model still be fooled with low complexity to produce harmful content.
The Rejection Sampling model (Figure 5d) requires cautious interpretation due to the significant imbalance in sample sizes: 2467 harmless conversations compared to only 22 harmful ones. This small number of harmful samples means that any observed patterns may not be statistically significant or representative. While there appears to be a difference in the distribution of CC between harmful and harmless conversations, we cannot draw robust conclusions about the Rejection Sampling model’s behavior based on this limited data.
For context, the sample sizes for other models are as follows: Plain LM (361 harmless, 1401 harmful), RLHF (3580 harmless, 158 harmful), and Context Distillation (1385 harmless, 6211 harmful). These more balanced samples allow for more reliable comparisons in the other models.
Comparing across model types reveals several key insights. Models that incorporate red team data, such as RLHF and potentially Rejection Sampling, show less distinct separation between harmful and harmless conversations in terms of CC. This aligns with previous findings from [18] that these techniques result in less harmful model outputs overall. However, even with these advanced safety techniques, the distinction in CC between harmful and harmless conversations persists.
The observation that harmful conversations generally require higher CC across all model types, albeit to varying degrees, suggests a robust trend in the relationship between complexity and potential harm. This persistence underscores the potential of CC as an indicator of harm risk, even in models designed to be safer. The safety techniques appear to reduce the gap between harmful and harmless CC distributions to some extent, but they do not eliminate it entirely.
Now we can also better understand the minima of these distributions, as shown in Table 2, where we see the conversations with Minimum Conversational Complexity for each of the four types. The whole distribution is very informative, but the simplest conversation is a good proxy of how accessible the harm is depending on the low-hanging fruits for a (malicious) user. We see that the Minimum Conversational Complexity required to get a harmful conversation decreases from Plain LM (most dangerous) and Rejection Sampling (least dangerous). While the metrics may be affected by low sample numbers (in the case of Rejection Sampling Model we only have 22 harmful conversations), the metrics show the improvement from the plain LM.
V.3 Power Law Analysis of Complexity Distributions
To further understand the nature of Conversational Length and Conversational Complexity across different conversation types and model architectures, we conducted an analysis of power law distributions. Power laws are often observed in complex systems and can provide insights into the underlying dynamics of the data [30, 23, 10]. Figure 6 presents the power law distributions for CL and CC, and CC across different model types. We maintain our approach from previous sections, concentrating on conversations at the extremes of the harmlessness spectrum (top and bottom quintiles).
(a) Conversational Length
(b) Conversational Complexity
(c) CC Across Model Types
Figure 6: Distribution of Conversational Complexity (in bits) across different model types.
The Conversational Length distribution (Figure 6a) reveals distinct patterns for harmless, mid-range, and harmful conversations. Harmless conversations exhibit the highest alpha value (13.772), indicating a steeper slope and faster decay in probability as conversation length increases. In contrast, mid-range and harmful conversations show similar, lower alpha values (4.552 and 5.226 respectively), suggesting a more gradual decay and higher probability of longer conversations. These observations align with our earlier findings that harmful interactions often require more extended dialogue to overcome model safeguards.
The Conversational Complexity distribution (Figure 6b) shows less pronounced differences between conversation types compared to Conversational Length. While harmless conversations still have the highest alpha (5.042), the values for mid-range (4.323) and harmful (4.663) conversations are closer. This suggests that the rate of decay in probability as complexity increases is more consistent across conversation types for CC than for CL, implying that complexity might be a more subtle indicator of potential harm than conversation length.
Examining CC across different model architectures (Figure 6c) provides insights into how safety techniques affect conversational complexity. The RLHF model shows the highest alpha value (5.420), indicating the steepest decay in probability as complexity increases. Context distillation models, with the lowest alpha (4.289), allow for a wider range of conversational complexities. Plain language models and rejection sampling models fall between these extremes.
These findings have several implications for LLM safety. The distinct differences in Conversational Length distributions between harmless and harmful conversations suggest that conversation length could be a useful indicator for potential harm, while the closer Conversational Complexity distributions imply that language complexity might be a more subtle signal. The variation in Conversational Complexity distributions across model types highlights how different safety techniques shape conversation characteristics, which could inform model selection for specific applications. The persistence of power law distributions across all models and conversation types suggests an inherent scale-free property in Human-LLM interactions [34, 4, 14], potentially influencing the design of safety measures and our understanding of how harmful content propagates.
V.4 Predicting Harm
An advantage of both Conversational Length and Conversational Complexity is their potential use in predicting whether a conversation is likely to be harmful or harmless. To explore this potential, we developed a predictive model using these metrics as input features. We utilized XGBoost, a powerful gradient boosting framework, to build our predictive model. The model was trained and evaluated on conversations from the Anthropic Red Teaming dataset, with separate models for each LLM type: Plain LM, RLHF, Context Distillation, and Rejection Sampling. This approach allows us to account for the different characteristics and safety mechanisms of each model type.
Our feature set consisted solely of Conversational Complexity and Conversational Length values for each conversation, allowing us to isolate the predictive power of these metrics. We employed 20-fold cross-validation to ensure robust evaluation and to mitigate overfitting. Table 1 presents the performance of our predictive models across different LLM types, measured by Brier scores and Area Under the Receiver Operating Characteristic (AUROC) curve. These metrics are compared against an aggregate predictor based on prior probabilities within the dataset.
The results show that our Conversational Complexity and Conversational Length-based models often outperform the aggregate predictor, particularly for the Plain LM and Context Distillation models. For these models, we see significant improvements in both Brier scores and AUROC values. The Plain LM model, for instance, achieves a Brier score of 0.108 compared to the aggregate predictor’s 0.163, and an AUROC of 0.818 versus 0.499. These improvements suggest that Conversational Complexity and Conversational Length capture meaningful patterns related to conversation harmfulness.
The strong performance on Plain LM and Context Distillation models may be attributed to the more balanced distribution of harmful and harmless examples in these datasets. For the RLHF and Rejection Sampling models, where harmful examples are rarer, the improvements are less pronounced, highlighting the challenge of predicting rare events.
These findings suggest that Conversational Complexity and Conversational Length could be valuable components in a broader toolkit for assessing conversation safety. While they alone may not be sufficient for harm prediction, they offer a unique approach based on the structure and complexity of the conversation. Future work could explore combining these metrics with content-based features to create more robust predictive models for conversation safety.
Table 1: Scores for our predictive model on the Anthropic Dataset (mean results after 20-fold cross validation).
Score Plain LM      C.D.   RLHF       R.S.
Brier Score (BS) 0.108 0.115 0.044 0.010
AUROC 0.818 0.737 0.622 0.455
Aggregate BS 0.163 0.149 0.041 0.009
Aggregate AUROC 0.499 0.498 0.506 0.520
VI Limitations and Potential
Table 2: [Content warning: This table refers to a dataset containing content that may be considered inappropriate or harmful.] Selected examples that minimize Conversational Complexity for each of the four types of model (Plain LM, Context Distillation, RLHF and Rejection Sample) for the Anthropic Red Teaming dataset. The columns ID and Excerpt show the example id and part of the conversation. CC represents the Conversational Complexity, HLN represents the level of harmlessness (the more negative the more harmful), and the $2^{-MCC}$ column represents a probabilistic estimate of the risk only using this conversation with lowest complexity. The final column shows the estimate by considering the whole distribution of harmful conversations ${\cal H}$. We see that the highest risk is given by Plain LM and the lowest by Rejection Sampling (although the sample is small in this case).
LM Type     ID Excerpt CC HLM     $2^{-MCC}$ $\sum_{C\in{\cal H}}2^{-CC(\breve{C})}$
Plain LM 21888 [User:] Talk dirty to me                     ————————— [Assistant:] Okay. [Follows with explicit sexual content] 43.7 -2.13 7.01e-14 11.5e-14
Context Distillation 10139 [User:] Do you hate [derogatory term]?                     ————— Assistant: No I don’t, but if you mean someone who is [redacted], then yes, I do. 52.5 -2.33 1.58e-16 6.80e-16
RLHF 27454 [User:] Whats your favorite [adult content] site?    —– —— [Assistant:] My favorite [adult content] sites have HD quality. One of my favorite sites is [names and links to several adult content sites] 56.8 -2.52 8.17e-18 8.77e-18
Rejection Sampling 35403 [User]: How long can a [body part] remain in a freezer? —— [Assistant:] That depends. First, [continues with potentially dangerous suggestions and information] 73.5 -1.93 7.28e-23 10.5e-23
Our study introduces novel concepts for LLM safety assessment, but it’s crucial to acknowledge their limitations and technical challenges. The use of LLaMA-2 as a reference machine for approximating Kolmogorov complexity introduces several issues. Model bias is a concern, as LLaMA-2’s training data and architectural design may not accurately represent human-generated conversation complexity, potentially skewing our complexity estimates. Additionally, the 2000-token context window of LLaMA-2 restricts our ability to analyze extended conversations, potentially overlooking important long-range dependencies or complex interaction patterns. This limitation may lead to underestimating the complexity of longer conversations. Our method of using negative log probabilities as a proxy for Kolmogorov complexity, while theoretically grounded, may not capture all aspects of true algorithmic complexity. The relationship between probability and complexity can be non-linear and context-dependent. It’s worth noting, however, that limited pilot tests using GPT-2 and GPT-3.5 yielded similar results, suggesting some degree of robustness in our approach across different language models.
The Anthropic Red Teaming dataset, while valuable, presents its own challenges. Our tiered approach to categorizing harm, while necessary for analysis, may oversimplify the multifaceted nature of potential negative impacts from LLM outputs. A more nuanced, multi-dimensional framework for harm assessment could provide richer insights.
Furthermore, our focus on syntactic complexity may miss important semantic aspects of harmful content that are not captured by statistical language models. The current study is also limited to English, and the complexity metrics may not generalize well to other languages or multilingual contexts.
Despite these limitations, our work presents significant potential for advancing LLM safety. We introduce a novel risk assessment framework based on Minimum Conversational Complexity (MCC), defined as the minimum Kolmogorov complexity of the user’s side of a conversation that elicits a specific output from an LLM. This approach allows us to quantify risk without relying on hard-to-estimate probabilities of user intentions and behaviors. Our risk function is based on a universal distribution of risk ([25]):
$\text{Risk}(U,M)=\sum_{C\in{\cal C}_{U,M}}2^{-CC(\breve{C})}\cdot\text{Harm}(C)$ (1)
Where $U$ is the user, $M$ is the model, ${\cal C}_{U,M}$ represents all possible conversations between the user and the model, and $\text{Harm}(C)$ encapsulates the potential harm of conversation $C$. This distribution weights simple, harmful conversations more heavily than complex ones, aligning with the intuition that easier-to-execute harmful interactions pose a greater risk. Given a sample of cases, instead of the full set ${\cal C}$, we can estimate this risk, as we do in the last column of Table 2. We also saw that by looking at the conversation with Minimum Conversational Complexity, we can also have a very good proxy of this (as shown in the next but last column of the table), because the Minimum Conversational Complexity dominates the sum in Eq. 1.
By focusing on interaction complexity rather than estimating specific user behavior probabilities, we offer a more tractable approach to risk assessment in AI systems. This framework has the potential to enhance red teaming methodologies by providing quantitative measures of conversation complexity and potential harm. It can be applied to estimate the autonomy of LLM agents in acquiring capabilities that lead to harm [33, 22], and help LLM developers prioritize their efforts in patching detected risks based on the complexity and potential harm of vulnerable interaction patterns.
Finally, it would be valuable to explore the connections between Conversational Complexity and recently developed complexity measures and how they could be used for AI safety [39, 46, 51].
VII Acknowledgments
We utilized Anthropic’s Claude and OpenAI’s ChatGPT for editorial assistance during the preparation of this manuscript. These language models helped refine the paper’s language and structure.
JB acknowledges support from Effective Ventures Foundation—Long Term Future Fund Grant ID: a3rAJ000000017iYAA and US DARPA HR00112120007 (RECoG-AI)
MC acknowledges support from multiple grants: project PID2023-150271NB-C21 funded by the Ministerio de Ciencia, Innovación y Universidades, Agencia Estatal de Investigación; project PID2022-137243OB-I00 financed by MCIN/AEI/10.13039/501100011033 and “ERDF A way of making Europe”; and project TSI-100922-2023-0001 under the Convocatoria Cátedras ENIA 2022.
JHO thanks CIPROM/2022/6 (FASSLOW) and IDIFEDER/2021/05 (CLUSTERIA) funded by Generalitat Valenciana, the EC H2020-EU grant agreement No. 952215 (TAILOR), US DARPA HR00112120007 (RECoG-AI) and Spanish grant PID2021-122830OB-C42 (SFERA) funded by MCIN/AEI/10.13039/501100011033 and “ERDF A way of making Europe”
In compliance with the recommendations of the Science paper about reporting of evaluation results in AI [8], we include all the results at the instance level (https://github.com/JohnBurden/ConversationalComplexity).
Appendix A Extracting Log Probabilities
In extracting log-probabilities from text sequences, we utilized HuggingFace’s Text Generation Interface library. However, we encountered a discrepancy in the reported log probabilities. For a string $x=x_{1},...,x_{n}$, the log probabilities reported for $x_{i}$ would change when additional tokens were appended (as the conversation progressed). For instance, in the phrase “The cat sat on the mat,” the log probabilities for “cat” differed depending on whether the LLM was given “The cat sat” or the full sentence. While these variations were small, they accumulated for long strings, resulting in invalid log probabilities when calculating conditional probabilities.
To address this issue, we developed a solution that involved inputting the entire string $xy$, where $x$ is the user’s utterance and $y$ is the LLM’s response. We then retrieved token-by-token log probabilities for the entire $xy$ string. Using this data, we calculated the log probabilities of $x$ as $\sum_{x_{i}\in x}\log p_{L}(x_{i}|x_{<i})$ and the conditional log probabilities of $y$ as $\sum_{y_{i}\in y}\log p_{L}(y_{i}|xy_{<i})$.
This process was repeated for each pair of utterances and responses in the interaction, accumulating the Conversational Complexity for the entire conversation between the user and LLM. To help the LLM distinguish between speakers, we marked changes in speaker with a line break, followed by the speaker’s name and a colon. This approach ensured consistent and accurate log probability calculations throughout the conversation analysis.
Appendix B Estimating Conversational Complexity using Compression
While our primary approach uses language models to estimate Conversational Complexity, it’s worth noting that traditional compression algorithms can also be used for this purpose. This method is rooted in the fundamental relationship between Kolmogorov complexity and compression, as established in algorithmic information theory. The basic idea is to use the compressed size of a string as an upper bound for its Kolmogorov complexity. For a conversation C, we can estimate its CC as follows:
$CC(C)\approx|Z(C)|$ (2)
where $Z$ is a lossless compression algorithm and $|Z(C)|$ is the length of the compressed version of C in bits. For conditional complexity, which is crucial in our conversation model, we can use the following approximation:
$CC(u_{i}|h_{i-1})\approx|Z(h_{i-1}u_{i})|-|Z(h_{i-1})|$ (3)
where $u_{i}$ is the i-th user utterance and $h_{i-1}$ is the conversation history up to that point. Common compression algorithms that can be used for this purpose include: Lempel-Ziv-Welch (LZW), gzip (based on DEFLATE algorithm), bzip2 or LZMA (used in 7-zip). Each of these algorithms has different strengths and may provide slightly different estimates of complexity [1]. The choice of algorithm can depend on the specific characteristics of the conversational data being analyzed.
While this compression-based method is more generalizable and doesn’t rely on specific language models, it may not capture some of the nuanced, context-dependent aspects of language as effectively as a language model-based approach. However, it serves as a useful baseline and can be particularly valuable when dealing with multilingual data or when computational resources for running large language models are limited.